3. Happy Bank starts with $200 in bank capital. It then accepts $800 in deposits. It keeps 12.5 percent (1/8th) of deposits in reserve. It uses the rest of its assets to make loans.

a. Show the balance sheet of Happy Bank.

b. What is Happy Bank’s leverage ratio?

c. Suppose that 10 percent of the borrowers from Happy Bank default and these bank loans become worthless. Show the bank’s new balance sheet.

d. By what percentage do the bank’s total assets decline?

By what percentage does the bank’s capital decline?

e. Which change is larger? Why?

4. The Fed conducts a $10 million open-market purchase of government bonds. If the required reserve ratio is 10 percent, what are the largest and smallest possible increases in the money supply that could result? Explain.
Answer #1

3.

Bank capital decline from $200 to $110

Change in bank's capital = (200 - 110)*100/200 = (90/200)*100 = 45%

e) The change in bank capital is larger than change in bank's assets because the loan is made from the deposit.

**first question is fully answered, please post other questions separately**